# The Initial Value Problem for Harmonic maps of Cohomogeneity One manifolds

### v1

@@ 4. Regular-singular systems of first order
In the present section we present a theory for regular-singular systems of first order. Many of the results are known, but scattered throughout the literature. The organization of the section is as follows: In the first two subsections we consider the regular-singular differential equation/system whose solutions are power functions/vector-valued generalizations of power functions. Basic phenomena are explained with the help of these examples. Afterwards, in Subsections 4.3-4.7 we study linear regular-singular system of first order. In particular, we discuss the monodromy map. Finally, in Subsections 4.8-4.10 we study non-linear regular-singular system of first order.

### v2

@@ 6. Regular-singular systems of first order
In the present section we present a theory for regular-singular systems of first order. Many of the results are known, but scattered throughout the literature. The organization of the section is as follows: In the first two subsections we consider the regular-singular differential equation/system whose solutions are power functions/vector-valued generalizations of power functions. Basic phenomena are explained with the help of these examples. Afterwards, in Subsections 6.3-6.7 we study linear regular-singular system of first order. In particular, we discuss the monodromy map. Finally, in Subsections 6.8-6.10 we study non-linear regular-singular system of first order.